# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ 6.1 Results under the inner product latent space model
Table 2: Node-level statistic distribution preservation results with $n=4000$ under the inner product latent space model, measured by the average Wasserstein distance over 100 replications. The numbers in parentheses are the corresponding standard errors.
Metric $d$ $\varepsilon$ $\rho=0.025$ $\rho=0.05$ $\rho=0.1$
Hat GRAND Laplace Hat GRAND Laplace Hat GRAND Laplace
Node Degree 3 1 0.033 (<0.001<0.001) 0.049 (0.002) 3.04 (0.001) 0.01 (<0.001<0.001) 0.03 (0.001) 2.34 (0.001) 0.004 (<0.001<0.001) 0.02 (0.001) 1.635 (0.001)
2 0.033 (<0.001<0.001) 0.042 (0.002) 2.956 (0.002) 0.01 (<0.001<0.001) 0.03 (0.001) 2.276 (0.001) 0.004 (<0.001<0.001) 0.019 (0.001) 1.588 (0.001)
5 0.033 (<0.001<0.001) 0.031 (0.001) 2.552 (0.003) 0.01 (<0.001<0.001) 0.023 (0.001) 1.949 (0.003) 0.004 (<0.001<0.001) 0.017 (0.001) 1.341 (0.003)
10 0.033 (<0.001<0.001) 0.035 (0.001) 1.787 (0.005) 0.01 (<0.001<0.001) 0.018 (0.001) 1.327 (0.005) 0.004 (<0.001<0.001) 0.016 (0.001) 0.895 (0.004)
6 1 0.04 (<0.001<0.001) 0.053 (0.002) 3.05 (0.001) 0.012 (<0.001<0.001) 0.034 (0.001) 2.348 (0.001) 0.004 (<0.001<0.001) 0.018 (0.001) 1.644 (0.001)
2 0.04 (<0.001<0.001) 0.048 (0.001) 3.031 (0.001) 0.012 (<0.001<0.001) 0.03 (0.001) 2.334 (0.001) 0.004 (<0.001<0.001) 0.017 (0.001) 1.634 (0.001)
5 0.04 (<0.001<0.001) 0.036 (0.001) 2.91 (0.001) 0.012 (<0.001<0.001) 0.026 (0.001) 2.238 (0.001) 0.004 (<0.001<0.001) 0.015 (0.001) 1.565 (0.001)
10 0.04 (<0.001<0.001) 0.041 (0.001) 2.555 (0.003) 0.012 (<0.001<0.001) 0.022 (0.001) 1.957 (0.003) 0.004 (<0.001<0.001) 0.014 (0.001) 1.358 (0.002)
V-shape 3 1 0.068 (0.001) 0.1 (0.004) 6.113 (0.002) 0.02 (<0.001<0.001) 0.061 (0.003) 4.696 (0.002) 0.008 (<0.001<0.001) 0.041 (0.002) 3.276 (0.002)
2 0.068 (0.001) 0.085 (0.003) 5.945 (0.003) 0.02 (<0.001<0.001) 0.06 (0.003) 4.568 (0.002) 0.008 (<0.001<0.001) 0.039 (0.002) 3.182 (0.002)
5 0.068 (0.001) 0.063 (0.002) 5.137 (0.007) 0.02 (<0.001<0.001) 0.047 (0.002) 3.913 (0.007) 0.008 (<0.001<0.001) 0.035 (0.002) 2.689 (0.005)
10 0.068 (0.001) 0.071 (0.002) 3.603 (0.011) 0.02 (<0.001<0.001) 0.036 (0.001) 2.666 (0.009) 0.008 (<0.001<0.001) 0.032 (0.002) 1.796 (0.007)
6 1 0.082 (<0.001<0.001) 0.108 (0.003) 6.132 (0.002) 0.024 (<0.001<0.001) 0.069 (0.003) 4.712 (0.002) 0.009 (<0.001<0.001) 0.037 (0.002) 3.295 (0.001)
2 0.082 (<0.001<0.001) 0.098 (0.003) 6.095 (0.002) 0.024 (<0.001<0.001) 0.061 (0.003) 4.682 (0.002) 0.009 (<0.001<0.001) 0.035 (0.002) 3.274 (0.001)
5 0.082 (<0.001<0.001) 0.073 (0.002) 5.852 (0.003) 0.024 (<0.001<0.001) 0.054 (0.002) 4.491 (0.003) 0.009 (<0.001<0.001) 0.03 (0.002) 3.136 (0.002)
10 0.082 (<0.001<0.001) 0.084 (0.002) 5.142 (0.006) 0.024 (<0.001<0.001) 0.045 (0.002) 3.929 (0.006) 0.009 (<0.001<0.001) 0.028 (0.001) 2.722 (0.004)
Triangle 3 1 0.024 (<0.001<0.001) 0.097 (0.005) 8.759 (0.004) 0.013 (<0.001<0.001) 0.081 (0.005) 6.768 (0.003) 0.008 (<0.001<0.001) 0.062 (0.004) 4.796 (0.003)
2 0.024 (<0.001<0.001) 0.088 (0.004) 8.587 (0.005) 0.013 (<0.001<0.001) 0.079 (0.004) 6.633 (0.003) 0.008 (<0.001<0.001) 0.06 (0.004) 4.695 (0.003)
5 0.024 (<0.001<0.001) 0.094 (0.004) 7.747 (0.007) 0.013 (<0.001<0.001) 0.066 (0.004) 5.933 (0.007) 0.008 (<0.001<0.001) 0.055 (0.003) 4.148 (0.006)
10 0.024 (<0.001<0.001) 0.116 (0.005) 6.052 (0.012) 0.013 (<0.001<0.001) 0.054 (0.003) 4.491 (0.011) 0.008 (<0.001<0.001) 0.052 (0.003) 3.05 (0.01)
6 1 0.024 (<0.001<0.001) 0.091 (0.004) 8.783 (0.004) 0.015 (<0.001<0.001) 0.073 (0.004) 6.765 (0.003) 0.01 (<0.001<0.001) 0.049 (0.003) 4.777 (0.002)
2 0.024 (<0.001<0.001) 0.1 (0.005) 8.746 (0.004) 0.015 (<0.001<0.001) 0.068 (0.004) 6.734 (0.003) 0.01 (<0.001<0.001) 0.043 (0.003) 4.754 (0.002)
5 0.024 (<0.001<0.001) 0.12 (0.005) 8.496 (0.004) 0.015 (<0.001<0.001) 0.064 (0.003) 6.531 (0.004) 0.01 (<0.001<0.001) 0.037 (0.002) 4.604 (0.002)
10 0.024 (<0.001<0.001) 0.169 (0.006) 7.749 (0.007) 0.015 (<0.001<0.001) 0.059 (0.003) 5.921 (0.007) 0.01 (<0.001<0.001) 0.037 (0.002) 4.139 (0.004)
Eigen Centrality 3 1 0.009 (0.001) 0.031 (0.002) 0.331 (0.005) 0.007 (<0.001<0.001) 0.028 (0.002) 0.257 (0.005) 0.007 (0.001) 0.028 (0.002) 0.164 (0.004)
2 0.009 (0.001) 0.031 (0.002) 0.292 (0.004) 0.007 (<0.001<0.001) 0.029 (0.002) 0.208 (0.004) 0.007 (0.001) 0.025 (0.002) 0.103 (0.002)
5 0.009 (0.001) 0.036 (0.002) 0.193 (0.004) 0.007 (<0.001<0.001) 0.028 (0.002) 0.115 (0.004) 0.007 (0.001) 0.023 (0.002) 0.044 (0.001)
10 0.009 (0.001) 0.037 (0.003) 0.063 (0.004) 0.007 (<0.001<0.001) 0.029 (0.002) 0.041 (0.002) 0.007 (0.001) 0.025 (0.002) 0.1 (0.003)
6 1 0.012 (0.001) 0.037 (0.003) 0.466 (0.007) 0.009 (0.001) 0.028 (0.002) 0.409 (0.007) 0.008 (0.001) 0.022 (0.002) 0.328 (0.004)
2 0.012 (0.001) 0.037 (0.003) 0.366 (0.006) 0.009 (0.001) 0.032 (0.002) 0.308 (0.006) 0.008 (0.001) 0.024 (0.002) 0.228 (0.005)
5 0.012 (0.001) 0.04 (0.003) 0.247 (0.005) 0.009 (0.001) 0.031 (0.002) 0.177 (0.005) 0.008 (0.001) 0.025 (0.002) 0.099 (0.003)
10 0.012 (0.001) 0.041 (0.003) 0.144 (0.005) 0.009 (0.001) 0.035 (0.002) 0.08 (0.004) 0.008 (0.001) 0.022 (0.002) 0.048 (0.001)
Harmonic Centrality 3 1 21.91 (0.115) 23.201 (0.742) 1036.187 (0.638) 2.046 (0.025) 4.445 (0.192) 884.79 (0.37) 0.738 (0.009) 3.886 (0.23) 784.981 (0.335)
2 21.91 (0.115) 20.024 (0.664) 976.434 (0.83) 2.046 (0.025) 4.301 (0.18) 836.624 (0.756) 0.738 (0.009) 3.722 (0.21) 747.901 (0.597)
5 21.91 (0.115) 10.857 (0.554) 732.359 (1.592) 2.046 (0.025) 3.298 (0.155) 621.66 (1.79) 0.738 (0.009) 3.381 (0.189) 571.464 (1.537)
10 21.91 (0.115) 6.869 (0.295) 417.292 (1.569) 2.046 (0.025) 2.374 (0.101) 324.391 (1.682) 0.738 (0.009) 3.109 (0.165) 317.947 (1.688)
6 1 23.68 (0.118) 24.162 (0.659) 1041.142 (0.376) 1.953 (0.018) 4.435 (0.177) 898.208 (0.168) 0.765 (0.009) 3.337 (0.176) 796.937 (0.139)
2 23.68 (0.118) 21.104 (0.646) 1026.115 (0.419) 1.953 (0.018) 3.904 (0.162) 885.7 (0.286) 0.765 (0.009) 3.116 (0.149) 787.725 (0.226)
5 23.68 (0.118) 12.597 (0.542) 933.849 (0.746) 1.953 (0.018) 3.362 (0.139) 808.732 (0.812) 0.765 (0.009) 2.683 (0.143) 729.535 (0.517)
10 23.68 (0.118) 7.174 (0.265) 713.174 (1.418) 1.953 (0.018) 2.785 (0.095) 618.093 (1.537) 0.765 (0.009) 2.524 (0.124) 575.932 (1.051)
6.2 Results under the RDPG model
For the RDPG model, we sample latent vectors $Z_{1},\ldots,Z_{N}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\text{Uniform}[0,1]^{d}$. The $Z_{i}$’s are then rescaled to ensure the desired network density. The other steps remain the same as in the previous experiment. Tables 3 and 4 present the experimental results under the RDPG model measured by the Wasserstein distance, corresponding to $n=2000$ and $n=4000$, respectively. The observed patterns align closely with those seen under the inner product models. Our proposed GRAND method clearly outperforms the naive Laplace mechanism across all settings. For larger privacy budgets $\varepsilon$, the preservation of network properties becomes comparable to that of the Hat mechanism. It is worth noting that, overall, model estimation under the RDPG framework becomes noticeably noisier (than the inner product model) in very sparse networks, resulting in less discernible trends with respect to $\varepsilon$ in many cases. This is likely due to the fact that the RDPG estimation is based on first-order information rather than the likelihood (Athreya et al., 2018). However, in denser scenarios, we still observe the expected improvement in property preservation as the privacy budget increases.
Table 3: Node-level statistic distribution preservation results with $n=2000$ under the RDPG model, measured by the average Wasserstein distance over 100 replications. The numbers in the parentheses are the corresponding standard errors.
Metric $d$ $\varepsilon$ $\rho=0.025$ $\rho=0.05$ $\rho=0.1$
Hat GRAND Laplace Hat GRAND Laplace Hat GRAND Laplace
Node Degree 3 1 0.032 (<0.001<0.001) 0.042 (0.001) 2.559 (0.004) 0.017 (<0.001<0.001) 0.031 (0.001) 2.037 (0.003) 0.009 (<0.001<0.001) 0.026 (0.001) 1.511 (0.002)
2 0.032 (<0.001<0.001) 0.043 (0.001) 1.715 (0.007) 0.017 (<0.001<0.001) 0.03 (0.001) 1.408 (0.005) 0.009 (<0.001<0.001) 0.027 (0.001) 1.126 (0.003)
5 0.032 (<0.001<0.001) 0.043 (0.001) 0.521 (0.003) 0.017 (<0.001<0.001) 0.028 (0.001) 0.554 (0.003) 0.009 (<0.001<0.001) 0.023 (0.001) 0.518 (0.003)
10 0.032 (<0.001<0.001) 0.047 (0.001) 0.47 (0.003) 0.017 (<0.001<0.001) 0.028 (0.001) 0.508 (0.002) 0.009 (<0.001<0.001) 0.021 (0.001) 0.484 (0.002)
6 1 0.036 (<0.001<0.001) 0.038 (0.001) 2.921 (0.001) 0.018 (<0.001<0.001) 0.025 (0.001) 2.269 (0.001) 0.01 (<0.001<0.001) 0.021 (0.001) 1.601 (0.001)
2 0.036 (<0.001<0.001) 0.039 (0.001) 2.631 (0.003) 0.018 (<0.001<0.001) 0.024 (0.001) 2.082 (0.002) 0.01 (<0.001<0.001) 0.019 (0.001) 1.486 (0.001)
5 0.036 (<0.001<0.001) 0.039 (0.001) 1.444 (0.006) 0.018 (<0.001<0.001) 0.023 (0.001) 1.179 (0.005) 0.01 (<0.001<0.001) 0.018 (0.001) 0.869 (0.004)
10 0.036 (<0.001<0.001) 0.04 (0.001) 0.516 (0.003) 0.018 (<0.001<0.001) 0.023 (0.001) 0.524 (0.003) 0.01 (<0.001<0.001) 0.017 (0.001) 0.567 (0.002)
V-shape 3 1 0.068 (0.001) 0.088 (0.003) 5.183 (0.008) 0.034 (<0.001<0.001) 0.065 (0.003) 4.106 (0.006) 0.019 (<0.001<0.001) 0.053 (0.002) 3.036 (0.003)
2 0.068 (0.001) 0.091 (0.003) 3.486 (0.014) 0.034 (<0.001<0.001) 0.062 (0.003) 2.841 (0.01) 0.019 (<0.001<0.001) 0.055 (0.002) 2.263 (0.006)
5 0.068 (0.001) 0.091 (0.003) 1.069 (0.006) 0.034 (<0.001<0.001) 0.058 (0.002) 1.129 (0.006) 0.019 (<0.001<0.001) 0.047 (0.002) 1.047 (0.006)
10 0.068 (0.001) 0.098 (0.003) 1.005 (0.006) 0.034 (<0.001<0.001) 0.057 (0.002) 1.069 (0.005) 0.019 (<0.001<0.001) 0.043 (0.002) 1 (0.004)
6 1 0.074 (0.001) 0.078 (0.002) 5.902 (0.002) 0.037 (<0.001<0.001) 0.051 (0.002) 4.567 (0.002) 0.019 (<0.001<0.001) 0.042 (0.002) 3.215 (0.002)
2 0.074 (0.001) 0.08 (0.002) 5.321 (0.005) 0.037 (<0.001<0.001) 0.048 (0.002) 4.191 (0.004) 0.019 (<0.001<0.001) 0.039 (0.002) 2.984 (0.003)
5 0.074 (0.001) 0.081 (0.002) 2.935 (0.013) 0.037 (<0.001<0.001) 0.047 (0.002) 2.378 (0.011) 0.019 (<0.001<0.001) 0.037 (0.002) 1.747 (0.008)
10 0.074 (0.001) 0.084 (0.002) 1.055 (0.007) 0.037 (<0.001<0.001) 0.048 (0.002) 1.063 (0.006) 0.019 (<0.001<0.001) 0.035 (0.002) 1.145 (0.005)
Triangle 3 1 0.075 (0.001) 0.095 (0.004) 8.209 (0.008) 0.037 (0.001) 0.078 (0.004) 6.53 (0.006) 0.019 (<0.001<0.001) 0.066 (0.003) 4.788 (0.004)
2 0.075 (0.001) 0.091 (0.004) 6.208 (0.017) 0.037 (0.001) 0.076 (0.004) 5.105 (0.012) 0.019 (<0.001<0.001) 0.067 (0.003) 3.944 (0.007)
5 0.075 (0.001) 0.085 (0.003) 2.366 (0.016) 0.037 (0.001) 0.066 (0.003) 2.135 (0.01) 0.019 (<0.001<0.001) 0.057 (0.003) 1.853 (0.008)
10 0.075 (0.001) 0.08 (0.003) 1.199 (0.008) 0.037 (0.001) 0.058 (0.003) 1.373 (0.008) 0.019 (<0.001<0.001) 0.051 (0.002) 1.371 (0.006)
6 1 0.127 (0.001) 0.1 (0.003) 8.981 (0.003) 0.052 (0.001) 0.067 (0.003) 7.005 (0.003) 0.023 (<0.001<0.001) 0.055 (0.003) 5.005 (0.003)
2 0.127 (0.001) 0.102 (0.003) 8.312 (0.006) 0.052 (0.001) 0.062 (0.003) 6.583 (0.005) 0.023 (<0.001<0.001) 0.05 (0.003) 4.753 (0.003)
5 0.127 (0.001) 0.092 (0.003) 5.332 (0.018) 0.052 (0.001) 0.058 (0.003) 4.442 (0.014) 0.023 (<0.001<0.001) 0.046 (0.002) 3.37 (0.01)
10 0.127 (0.001) 0.085 (0.002) 2.314 (0.017) 0.052 (0.001) 0.056 (0.002) 2.049 (0.011) 0.023 (<0.001<0.001) 0.042 (0.002) 1.908 (0.008)
Eigen Centrality 3 1 0.051 (0.002) 0.049 (0.002) 0.115 (0.002) 0.031 (0.002) 0.033 (0.002) 0.14 (0.002) 0.019 (0.001) 0.023 (0.002) 0.137 (0.001)
2 0.051 (0.002) 0.05 (0.003) 0.187 (0.002) 0.031 (0.002) 0.033 (0.002) 0.18 (0.001) 0.019 (0.001) 0.021 (0.001) 0.164 (0.001)
5 0.051 (0.002) 0.047 (0.003) 0.299 (0.003) 0.031 (0.002) 0.034 (0.002) 0.284 (0.002) 0.019 (0.001) 0.02 (0.001) 0.209 (0.002)
10 0.051 (0.002) 0.052 (0.003) 0.263 (0.003) 0.031 (0.002) 0.032 (0.002) 0.286 (0.003) 0.019 (0.001) 0.022 (0.002) 0.268 (0.002)
6 1 0.075 (0.003) 0.054 (0.003) 0.303 (0.003) 0.042 (0.002) 0.036 (0.002) 0.272 (0.004) 0.021 (0.001) 0.029 (0.002) 0.237 (0.004)
2 0.075 (0.003) 0.053 (0.003) 0.08 (0.002) 0.042 (0.002) 0.036 (0.002) 0.092 (0.002) 0.021 (0.001) 0.027 (0.002) 0.12 (0.002)
5 0.075 (0.003) 0.054 (0.003) 0.264 (0.002) 0.042 (0.002) 0.036 (0.002) 0.233 (0.002) 0.021 (0.001) 0.028 (0.002) 0.202 (0.001)
10 0.075 (0.003) 0.05 (0.003) 0.337 (0.003) 0.042 (0.002) 0.037 (0.002) 0.339 (0.003) 0.021 (0.001) 0.026 (0.002) 0.278 (0.002)
Harmonic Centrality 3 1 3.548 (0.059) 6.465 (0.34) 394.221 (1.054) 1.628 (0.03) 3.133 (0.157) 332.865 (0.911) 0.83 (0.012) 2.388 (0.111) 331.579 (0.533)
2 3.548 (0.059) 6.846 (0.336) 207.534 (1.267) 1.628 (0.03) 2.99 (0.163) 169.035 (1.21) 0.83 (0.012) 2.484 (0.115) 213.128 (0.867)
5 3.548 (0.059) 6.853 (0.328) 44.747 (0.388) 1.628 (0.03) 2.866 (0.129) 76.293 (0.556) 0.83 (0.012) 2.103 (0.093) 92.078 (0.485)
10 3.548 (0.059) 7.361 (0.323) 52.71 (0.551) 1.628 (0.03) 2.891 (0.145) 81.309 (0.75) 0.83 (0.012) 1.887 (0.081) 67.14 (0.378)
6 1 4.03 (0.037) 5.486 (0.206) 526.587 (0.342) 1.563 (0.02) 2.144 (0.087) 427.069 (0.2) 0.877 (0.012) 1.978 (0.099) 382.348 (0.19)
2 4.03 (0.037) 5.613 (0.229) 419.844 (0.834) 1.563 (0.02) 2.027 (0.096) 352.837 (0.646) 0.877 (0.012) 1.805 (0.08) 335.148 (0.505)
5 4.03 (0.037) 5.68 (0.21) 182.552 (0.893) 1.563 (0.02) 2.018 (0.083) 136.861 (0.998) 0.877 (0.012) 1.687 (0.079) 165.261 (0.983)
10 4.03 (0.037) 6.044 (0.214) 56.276 (0.413) 1.563 (0.02) 2.025 (0.081) 58.349 (0.389) 0.877 (0.012) 1.603 (0.077) 88.231 (0.481)
@@ 7 Examples of real-world data sets
In this section, we demonstrate our method on privatizing real-world social network data sets. Similar to the simulation experiments, we focus on demonstrating how well the released network preserves the network properties, while fixing the privacy budget. In addition to the Hat method and the Laplace mechanism, we also use the true local network statistics of the original network as an addition benchmark.
Caltech Facebook network

### v3

@@ 6 Simulation Experiments
    Figure 3: Comparison of node-level statistic distribution preservation for $\rho=0.05$ with $n=4000$ and $d=6$ under the inner product latent space model (LSM) and random dot product graph (RDPG) model. Values are the average Wasserstein distance over 100 replications. The average distances are more than ten times larger than their corresponding standard errors; therefore, standard errors are omitted for brevity.
    Metric $\varepsilon$ LSM RDPG
    Hat GRAND Laplace Hat GRAND Laplace
    Node Degree 1 0.012 0.034 2.348 0.010 0.016 2.259
    2 0.012 0.030 2.334 0.010 0.015 2.038
    5 0.012 0.026 2.238 0.010 0.015 1.074
    10 0.012 0.022 1.957 0.010 0.015 0.562
    V-shape Count 1 0.024 0.069 4.712 0.020 0.032 4.532
    2 0.024 0.061 4.682 0.020 0.031 4.090
    5 0.024 0.054 4.491 0.020 0.031 2.158
    10 0.024 0.045 3.929 0.020 0.030 1.134
    Triangle Count 1 0.015 0.073 6.765 0.023 0.039 6.993
    2 0.015 0.068 6.734 0.023 0.038 6.501
    5 0.015 0.064 6.531 0.023 0.036 4.226
    10 0.015 0.059 5.921 0.023 0.034 2.036
    Eigen Centrality 1 0.009 0.028 0.409 0.022 0.024 0.234
    2 0.009 0.032 0.308 0.022 0.025 0.117
    5 0.009 0.031 0.177 0.022 0.028 0.255
    10 0.009 0.035 0.080 0.022 0.023 0.355
    Harmonic Centrality 1 1.953 4.435 898.208 1.042 1.670 834.276
    2 1.953 3.904 885.700 1.042 1.617 666.549
    5 1.953 3.362 808.732 1.042 1.596 239.014
    10 1.953 2.785 618.093 1.042 1.528 120.293
  • Comparing our method with the Hat method, we can see that it indeed yields worse preservation. This is also expected since it reflects the cost of achieving privacy. Noticeably, for weak privacy protection, such as $\varepsilon=10$, the performance of our method becomes comparable to the non-private Hat method. Recall that the Hat method is based on an estimation of a size-$m$ network. This verifies our theory that the node-wise estimation basically gives the accuracy as a standard estimation of size-$m$ network.
  • Comparing our method with the Laplace mechanism, where both methods satisfy the same strictness of node DP, it is evident that our method achieves much better preservation of network properties. The Laplace mechanism blindly introduces the noises but fails to preserve the network structures.
  • The patterns are consistent across configurations (see Appendix I), demonstrating the effectiveness of GRAND across a wide range of scenarios.
@@ 7 Examples of real-world data sets
We now demonstrate how well the released network preserves the network properties, fixing $\varepsilon$, on real-world networks. In addition to the Hat method and the Laplace mechanism, we also include the network statistics of the original network (True) in comparisons.